Consider the transportation table below. The solution displayed was obtained by performing some iterations of the transportation method on this problem.

                                                                                                                                

Consider the transportation table below. The solution displayed was obtained by performing some iterations of the transportation method on this problem.

                                                                                                                                

Copy to Clipboard +

Destination

Source

DenverDenver

YumaYuma

MiamiMiami

Supply

HoustonHouston

​$22

​$88

​$11

2020

2020

St. LouisSt. Louis

​$44

​$55

​$66

3030

1010

2020

ChicagoChicago

​$66

​$33

​$22

4040

1010

3030

Demand

3030

3030

3030

The total cost of the current solution provided​ =

​$270 270.

Using stepping stone​ method, to perform an improvement​ iteration, the improvement index for the unused cell will​ be:

HoustonHoustonminus−YumaYumaequals=

5 5

HoustonHoustonminus−MiamiMiamiequals=

negative 1−1

St. LouisSt. Louisminus−MiamiMiamiequals=

nothing

Destination

Source

Denver

Yuma

Miami

Supply

Houston

​$2

​$8

​$1

20

20

St. Louis

​$4

​$5

​$6

30

10

20

Chicago

​$6

​$3

​$2

40

10

30

Demand

30

30

30

The total cost of the current solution provided​ =

​$270.

Using stepping stone​ method, to perform an improvement​ iteration, the improvement index for the unused cell will​ be:

Houston - Yuma = 5

Houston - Miami = -1

St. Louis - Miami= ????
Answer #1

The total cost of the current solution provided = 20*2+10*4+20*5+10*3+30*2 = $ 270

Improvement index of empty cells is calculated by alternatively adding and subtracting the cost of cells which form a closed loop the empty cell at one of its corner and non-empty cells at all the other corners.

Closed loop for empty cell Houston-Yuma is formed by cells (Houston-Yuma) - (St.Louis-Yume) + (St.Louis-Denver) - (Houston-Denver)

Houston-Yuma = 8-5+4-2 = 5

Closed loop for empty cell Houston-Miami is formed by cells (Houston-Yuma) - (Chicago-Miami) + (Chicago-Yuma) - (St.Louis-Yume) + (St.Louis-Denver) - (Houston-Denver)

Houston-Miami = 1-2+3-5+4-2 = -1

Closed loop for empty cell St.Louis-Miami is formed by cells (St.Louis-Miami) - (Chicago-Miami) + (Chicago-Yuma) - (St.Louis-Yuma)

St.Louis-Miami = 6-2+3-5 = 2

Houston - Yuma = 5

Houston - Miami = -1

St. Louis - Miami= 2